NoGrowth Corporation currently pays a dividend of $0.44 per quarter, and it will continue to pay this dividend forever. What is the price per share of NoGrowth stock if the firms equity cost of capital is 13.7%? The stock price is $ (Round to the nearest cent.)
Answer #1

Answer:

Dividend D=$0.44 per quarter

Cost of capital=13.7% per annum

Per quarter cost of capital r=13.7%/4=3.425%

So price of share=D/r=0.44/3.425%=$12.85 per share